OpenStudy (anonymous):

3z + x = -4 y + 5z = 1 6x + 2z = 3 -x -y -z = 4 will the above form a linear system?

OpenStudy (anonymous):

Are you asking if the system is consistent?

OpenStudy (anonymous):

The answer is no, the system is inconsistent. You can show this with row reduction.

OpenStudy (anonymous):

yes you can get a linear equation from the above equations.Put the values of x and y in the last equation and you will get the value of z which equals to1/7 and then y=2/7 while the value of x=-31/7.

OpenStudy (anonymous):

This is wrong, substitute into 6x+2z=3 you will see it doesnt work.

OpenStudy (anonymous):

\[6\frac{-31}{7}+2\frac{1}{7} =-\frac{184}{7} \]